OpenStudy (anonymous):

Which expression is equivalent to this one? –5(6 + 11) A. (–5) · 6 + 11 B. (–5) + 6 + 11 C. (–5) · 6 · 11 D. (–5) · 6 + (–5) · 11

OpenStudy (timaashorty):

–5(6 + 11) = -85 \[A.(–5) · 6 + 11 = -19\] \[B.(–5) + 6 + 11 = 12 \] \[C.(–5) · 6 · 11 = -330 \] \[D.(–5) · 6 + (–5) · 11 = -85 \] \[\large Therefore,\ D. \ is \ equivalent. \]

OpenStudy (anonymous):

thanks! :)

OpenStudy (timaashorty):

You're welcome (: